OpenStudy (rissyroo118):

Question 8 of 20In the function y = 0.25x + 3.5, y represents the cost of a gallon of milk after a certain number of years, x. How much does the cost of milk increase every year? (Points : 5) $0.07 $0.25 $0.60 $14

OpenStudy (gorv):

for year 1 that is x=1 cost=y=0.25+3.5==3.75 for x=2 cost=y=0.25*2+3.5=0.50+3.5=4 increase per year=4-3.75

OpenStudy (rissyroo118):

$0.25 right?

OpenStudy (gorv):

yeppp right

OpenStudy (rissyroo118):

I can buy 56 diapers at Babies-R-Us for $28, and I can buy 112 diapers online for $52. Which is the better deal, and how much will you save per diaper? (Points : 5) Babies R Us by $0.04 Online by $0.04 Babies R Us by $0.15 Online by $0.15